Counting Principle:

Keeps the phone numbers of her seven best friends (4 women and 3 men) in her digital phone. In how many ways can she list them if all of the women are listed first?


I need to know how to work this problem please


1
Expert's answer
2021-10-29T04:36:18-0400

the number of ways can she list them if all of the women are listed first

"{4!}\times {3!}\\=(4 \times 3 \times 2 \times 1 ) \times ( 3 \times 2 \times 1)\\=24 \times 6\\=144"